The runner practices by running around a rectangular field 375 m by 250 m. How many times must he go around the field to run a distance of 5 km?
1
Expert's answer
2010-07-21T10:39:31-0400
We can find the perimeter of the field: P=(a+b)*2; a=375 m; b=250 m; P=(375+250)*2=1250 m - distance per lap. n=5000m/1250m=4 laps. So, the runner must go 4 times around the field to run a distance of 5 km.